# Adapting Noise to Data: Generative Flows from 1D Processes

### v3

@@ 3 From One-Dimensional to Multi-Dimensional Flows
These three examples highlight the adaptability of the framework and the different choices that can be made. For simplicity our approach for learning the latent distribution will only consider processes defined as a deterministic scaling of a fixed random variable. Therefore, in the last part of this section we also consider this case and how it fits into our framework.
3.1 Wiener Process and Diffusion Equation
@@
However, its $L_{2}$-norm fulfills $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}^{2}=\frac{d}{4t}$, and is therefore not integrable over time, i.e. $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}\notin L_{2}(0,1)$. In practice, instability issues caused by this explosion at times close to the target need to be avoided by e.g. time truncations, see e.g. [soft2021]. For a heuristic analysis also including drift-diffusion flows, we refer to [P2022]. Note that in the case of diffusion, there is no significant distinction between the uni- and multivariate setting. Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Wiener process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
Figure 3: A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the Wiener process.
3.2 Kac Process and Damped Wave Equation
Figure 4: Samples (1M) from the distribution of the two-dimensional Kac process $\mathbf{K}_{1}$ for $(a,c)=(9,3)$. We can clearly observe the atomic component of its distribution derived in  (18).
The Kac process was recently used in generative modeling by some of the authors in [DCFS2025]. It originates from a discrete random walk, which starts in 0 and moves with velocity parameter $c>0$ in one direction until it reverses its direction with probability $a\Delta_{t}$, $a>0$, see [KAC1974]. A continuous-time analogue is given by the Kac process which is defined using the homogeneous Poisson point process $N_{t}$ with rate $a$, i.e. i) $N_{0}=0$; ii) the increments of $N_{t}$ are independent, iii) $N_{t}-N_{s}\sim{\rm Poi}\big(a(t-s)\big)$ for all $0\leq s<t$. Now the Kac process starting in 0 is given by
$K_{t}:={\rm B}_{\frac{1}{2}}\,c\int_{0}^{t}(-1)^{N_{s}}\mathop{}\!\mathrm{d}s,$ (17)
@@ 3.2 Kac Process and Damped Wave Equation
By the following proposition, the Kac process is related to the damped wave equation, also known as telegrapher’s equation, and its probability distribution admits a computable vector field such that the continuity equation is fulfilled. For a proof we refer to our paper [DCFS2025].
Proposition 4.
@@
Interestingly, the damped wave equation (20) is closely related to the diffusion equation via Kac’ insertion method. In particular, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". Note that the diffusion-related concept of particles traveling with infinite speed violates Einstein’s laws of relativity and has therefore found resistance in the physics community [C1958, C1963, V1958, TL2016]. Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Kac process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
Figure 5: A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the Kac process with $(a,c)=(9,3)$.
3.3 Uniform Process and Gradient Flow of MMD Functional
Wasserstein gradient flows are special absolutely continuous measure flows, whose velocity fields are negative Wasserstein (sub-)gradients of functionals $\mathcal{F}_{\nu}$ on $\mathcal{P}_{2}(\mathbb{R}^{d})$ with the unique minimizer, i.e., $v_{t}\in-\partial\mathcal{F}_{\nu}(\mu_{t})$. The gradient descent flow should reach this minimizer as $t\to\infty$.
In this context, the MMD functional with the non-smooth negative distance kernel K(x,y)=|xy|K(x,y)=-|x-y| given by
$\mathcal{F}_{\nu}(\mu)=\text{MMD}^{2}_{K}(\mu,\nu)$ $\coloneqq-\frac{1}{2}\int_{\mathbb{R}^{2}}|x-y|\mathop{}\!\mathrm{d}\left(\mu(x)-\nu(x)\right)\mathop{}\!\mathrm{d}\left(\mu(y)-\nu(y)\right),$ (27)
stands out for its flexible flow behavior between distributions of different support [HGBS2022]. We have have the following proposition which proof can be found in Appendix A. Note that the proof is already based on so-called quantile function, which we consider in the next section.
Proposition 5.
@@ 3.3 Uniform Process and Gradient Flow of MMD Functional
It holds $\|v_{t}\|^{2}_{L_{2}(\mathbb{R},\mu_{t})}=\frac{2b}{3}\exp(-\frac{2t}{b})$, and hence, $\|v_{t}\|_{L_{2}(\mathbb{R},\mu_{t})}\in L_{2}(0,1)$. A corresponding (stochastic) process $(U_{t})_{t}$ is given by $U_{t}\coloneqq b\left(1-\exp\left(-\frac{t}{b}\right)\right)U$, where $U\sim\mathcal{U}[-1,1]$, such that $U_{t}\sim\mu_{t}$.
Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the process induced by the MMD gradient flow as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t$.
Figure 6: A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the MMD gradient flow.
3.4 Scaled Latent Distributions

### v4

@@ Appendix A Examples of One-Dimensional Flows
A.1 Wiener Process and Diffusion Equation
Figure 8: A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the Wiener process. As described in Section 2.4 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
@@ A.1 Wiener Process and Diffusion Equation
However, its $L_{2}$-norm fulfills $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}^{2}=\frac{d}{4t}$, and is therefore not integrable over time, i.e. $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}\notin L_{2}(0,1)$. In practice, instability issues caused by this explosion at times close to the target need to be avoided by e.g. time truncations, see e.g. (Kim et al., 2022). For a heuristic analysis also including drift-diffusion flows, we refer to (Pidstrigach, 2022). Note that in the case of diffusion, there is no significant distinction between the uni- and multivariate setting.
A.2 Kac Process and Damped Wave Equation
Figure 9: A generated trajectory from a flow matching model trained using the conditional density and velocity given by the Kac process with $(a,c)=(9,3)$. As described in Section 2.4 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
The Kac process (Kac, 1974), also known as persistent random walk, originates from a discrete random walk, which starts in 0 and moves with velocity parameter $c>0$ in one direction until it reverses its direction with probability $a\Delta_{t}$, $a>0$. A continuous-time analogue is given by the Kac process which is defined using the homogeneous Poisson point process $N_{t}$ with rate $a$, i.e. i) $N_{0}=0$; ii) the increments of $N_{t}$ are independent, iii) $N_{t}-N_{s}\sim{\rm Poi}\big(a(t-s)\big)$ for all $0\leq s<t$. Now the Kac process starting in 0 is given by
$K_{t}:={\rm B}_{\frac{1}{2}}\,c\int_{0}^{t}(-1)^{N_{s}}\,\mathrm{d}s,$ (30)
@@ A.2 Kac Process and Damped Wave Equation
By the following proposition, the Kac process is related to the damped wave equation, also known as telegrapher’s equation, and its probability distribution admits a computable vector field such that the continuity equation is fulfilled. For a proof we refer to (Duong et al., 2025).
Proposition A.1.
@@
Interestingly, the damped wave equation (33) is closely related to the diffusion equation via Kac’ insertion method. It is based on the following theorem, whose proof based on semigroup theory can be found in (Griego and Hersh, 1971), see also (Janssen, 1990; Kac, 1974).
Theorem A.2.
For any initial function $f_{0}\in H^{2}(\mathbb{R}^{d})$, $d\geq 1$, let $w_{c}(t,x)$ be the solution of the undamped wave equation with velocity $c>0$ given by
$\partial_{tt}w(t,x)$ $=c^{2}\Delta w(t,x),\quad x\in\mathbb{R}^{d},~t>0,$
$w(0,x)$ $=f_{0}(x),\qquad\partial_{t}w(0,x)=0.$ (40)
Then, the functions defined by
$h(t,x):=\mathbb{E}\left[w_{1}\left(\sigma W_{t},x\right)\right],\quad\text{resp.}\quad u(t,x):=\mathbb{E}\big[w_{c}(c^{-1}S_{t},x)\big]$ (41)
solve the diffusion equation
$\partial_{t}h(t,x)$ $=\frac{\sigma^{2}}{2}\Delta h(t,x),\quad x\in\mathbb{R}^{d},~t>0,$ (42)
$h(0,x)$ $=f_{0}(x),$ (43)
resp. the multi-dimensional damped wave equation
$\partial_{tt}u(t,x)+2a\,\partial_{t}u(t,x)$ $=c^{2}\Delta u(t,x),\quad x\in\mathbb{R}^{d},~t>0,$
$u(0,x)$ $=f_{0}(x),\qquad\partial_{t}u(0,x)=0.$ (44)
As a consequence, it is not hard to show the following corollary, see (Duong et al., 2025).
Corollary A.3.
For any $t\geq 0$, the solution $u^{a,c}(t,\cdot)$ of the damped wave equation (A.2) converges to the solution $h(t,\cdot)$ of the diffusion equation for $a,c\to\infty$ with fixed $\sigma^{2}=\frac{c^{2}}{a}$.
In other words, diffusion can be seen as ”an infinitely $a$-damped wave with infinite propagation speed $c$. Note that the diffusion-related concept of particles traveling with infinite speed violates Einstein’s laws of relativity and has therefore found resistance in the physics community (Cattaneo, 1958; Chester, 1963; Vernotte, 1958; Tautz and Lerche, 2016).
We also like to stress that in multiple dimensions, the damped wave equation (33) is no longer mass-conserving as in 1D (Tautz and Lerche, 2016), and hence eludes a characterization via stochastic processes. Figure 9 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Kac process as our noising process. As described in Section 2.4 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
A.3 Uniform Process and MMD Gradient Flow
Figure 10: A generated trajectory from a flow matching model trained using the conditional density and velocity given by the MMD gradient flow. As described in Section 2.4 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t$.
Wasserstein gradient flows are special absolutely continuous measure flows whose velocity fields are negative Wasserstein (sub-)gradients of functionals $\mathcal{F}_{\nu}$ on $\mathcal{P}_{2}(\mathbb{R}^{d})$ with the unique minimizer $\nu$. The gradient descent flow should reach this minimizer as $t\to\infty$. In this context, the MMD functional with the non-smooth negative distance kernel K(x,y)=|xy|{K(x,y)=-|x-y|} given by
$\mathcal{F}_{\nu}(\mu)=\text{MMD}^{2}_{K}(\mu,\nu)$ $\coloneqq-\frac{1}{2}\int_{\mathbb{R}^{2}}|x-y|\,\mathrm{d}\left(\mu(x)-\nu(x)\right)\,\mathrm{d}\left(\mu(y)-\nu(y)\right),$ (45)
stands out for its flexible flow behavior between distributions of different support (Hertrich et al., 2024). In 1D, its Wasserstein gradient flow $\mu_{t}$ can be equivalently described by the flow of its quantile functions $Q_{\mu_{t}}$ with respect to an associated functional on $L_{2}(0,1)$. Note that the MMD functional (45) loses its convexity (along generalized geodesics) in multiple dimensions (Hertrich et al., 2024), and the general existence of their Wasserstein gradient flows is unclear in the multivariate case. This yields another reason to work in 1D, where we have have the following proposition.
Proposition A.4.
@@ A.3 Uniform Process and MMD Gradient Flow
It holds $\|v_{t}\|^{2}_{L_{2}(\mathbb{R},\mu_{t})}=\frac{2b}{3}\exp(-\frac{2t}{b})$, and hence, $\|v_{t}\|_{L_{2}(\mathbb{R},\mu_{t})}\in L_{2}(0,1)$. A corresponding (stochastic) process $(U_{t})_{t}$ is given by $U_{t}\coloneqq b\left(1-\exp\left(-\frac{t}{b}\right)\right)U$, where $U\sim\mathcal{U}[-1,1]$, such that ${\rm Law}(U_{t})=\mu_{t}$.
We prove the proposition more general for $\nu=\mathcal{U}[a,b]$ and a flow starting in $x_{0}\in[a,b]$, i.e. we show
$\mu_{t}=\operatorname{\mathcal{U}}\left[a+(x_{0}-a)\exp\left(-r(t)\right),b-(b-x_{0})\exp\left(-r(t)\right)\right],\qquad t>0$ (48)
with $r(t)\coloneqq\frac{2t}{b-a}$ and
$v_{t}(x)=\frac{2}{b-a}\left(\frac{x-x_{0}}{\exp(r(t))-1}\right).$ (49)
To this end, we need the relation between measures in $\mathcal{P}_{2}(\mathbb{R})$ and cumulative distribution functions, see (15). For $\nu=\mathcal{U}[a,b]$, we have that
and $Q_{\nu}(s)=a(1-s)+bs$. In (Hertrich et al., 2024) it was shown that the functional $F_{\nu}\colon L_{2}(0,1)\to\mathbb{R}$ defined by
$F_{\nu}(u)\coloneqq\int_{0}^{1}\left((1-2s)\big(u(s)+Q_{\nu}(s)\big)+\int_{0}^{1}|u(s)-Q_{\nu}(t)|\,\mathrm{d}{t}\right)\,\mathrm{d}{s}$ (50)
fulfills $\mathcal{F}_{\nu}(\mu)=F_{\nu}(Q_{\mu})$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$. Moreover, we have the following equivalent characterization of Wasserstein gradient flows of $\operatorname{\mathcal{F}}_{\nu}$, which can be found in Theorem 4.5 (Duong et al., 2024).
Theorem A.5.
$v_{t}^{*}\circ g(t)\in-\partial F_{\nu}(g(t))\quad\text{and }\quad v_{t}^{*}\in-\partial\operatorname{\mathcal{F}}_{\nu}(\gamma_{t}).$ (52)
[…]
see Lemma 4.3 (Duong et al., 2024).
Proof of Proposition A.4.
We want to apply Theorem A.5 to $(\mu_{t},v_{t})$ in (48) and (49). The uniform distribution in (48) has the quantile function
$-\nabla F_{\nu}(Q_{\mu_{t}})(s)$ $=2s-2r_{\nu}(Q_{\mu_{t}}(s))$
$=2\left(s-\frac{x_{0}-a}{b-a}\right)\exp\left(-r(t)\right).$
On the other hand, it holds
By Theorem A.5, $(\mu_{t})$ is the unique Wasserstein gradient flow of $\operatorname{\mathcal{F}}_{\nu}$ starting in $\delta_{0}$.
$=\frac{2}{b-a}\left(\frac{y-x_{0}}{\exp\left(r(t)\right)-1}\right)$
$\|v_{t}\|_{L^{2}(\mathbb{R},\mu_{t})}^{2}$ $=\int\limits_{a+(x_{0}-a)\exp\left(-\frac{2t}{b-a}\right)}^{b-(b-x_{0})\exp\left(-\frac{2t}{b-a}\right)}\frac{4(x-x_{0})^{2}}{(b-a)^{2}\left(\exp\left(\frac{2t}{b-a}\right)-1\right)^{2}}\frac{1}{(b-a)\left(1-\exp\left(-\frac{2t}{b-a}\right)\right)}\,\mathrm{d}{x}$ (54)
$=\frac{4}{(b-a)^{3}\left(\exp\left(\frac{2t}{b-a}\right)-1\right)^{2}\left(1-\exp\left(-\frac{2t}{b-a}\right)\right)}\int\limits_{a+(x_{0}-a)\exp\left(-\frac{2t}{b-a}\right)}^{b-(b-x_{0})\exp\left(-\frac{2t}{b-a}\right)}(x-x_{0})^{2}\,\mathrm{d}{x}$ (55)
and the proof is finished. ∎
A.4 Scaled Latent Distributions
@@
Also note that if $Z=Q(U)$ for a quantile function $Q:(0,1)\to\mathbb{R}$ and a random variable $U\sim\mathcal{U}([0,1])$, we have
$\mathbb{E}\!\big[\|Z\|^{2}\big]=\int_{0}^{1}|Q(u)|^{2}\,\mathrm{d}u,$ (63)
i.e. the second moment of $Z$ is exactly given by the $L_{2}$-norm of its quantile. Hence, explosions of the velocity’s norm $\int_{0}^{1}\|v_{t}\|_{L_{2}(\mathbb{R},\mu_{t})}^{2}\,{\rm d}t$ can be directly controlled by the derivative of the time schedule $g$, and the size of the quantile function $Q$ of the latent variable.